# Nautilus: Locality-aware Autoencoder for Scalable Mesh Generation

### v4

@@ Appendix B Extensive Experiments and Analysis
Our experiments in the main paper comprehensively evaluate the effectiveness of Nautilus. In this section, we provide supplementary analysis and discussion to explore behind its efficacy:
@@
  • Sec.B.3: We analyze the modeling of local dependency by visualizing the attention maps in the transformer decoder.
  • Sec.B.4: We discuss the influence of sequence length on the model’s generation capability.
Figure 11. Ablated comparison of generated results between the representation of AMT and our Nautilus Shell representations. To independently analyze the influence of local dependency preservation, both groups are trained on the same dataset, and AMT is improved by incorporating the same coordinate compression as Nautilus to achieve a comparable number of faces, denoted as AMT Comp.
B.1. Quantitative Ablation Study
In our main paper, we conduct a qualitative ablation study on point-cloud-conditioned generation. Here, we extend the analysis to a quantitative evaluation on the same high-quality test set of 500 samples used in our main paper. In Table 3, we compare four configurations: using only the Coordinate Compression (Comp. only), using only the Nautilus Shell Representation (Rep. only), using our Full Tokenization but without the local encoder, and our Full Version.
@@ B.1. Quantitative Ablation Study
Table 3. We conduct quantitative ablation comparison on four specific configurations: using only the Coordinate Compression (Comp. only), using only the Nautilus Shell Representation (Rep. only), using our Full Tokenization but without the local encoder, and our Full Version.
@@
B.2. Ablated Comparison with AMT Representation
@@
Specifically, we apply the same coordinate compression to the AMT representation and train it on the same dataset as Nautilus, enabling it to model more than 5,000 faces. In Fig. 11, we denote the improved AMT version as AMT Comp., comparing it with our full version on challenging samples with complex topology and rich geometric details. Despite achieving a compression ratio comparable to that of Nautilus, AMT Comp. still exhibits severe manifold defects in its generation results, particularly isolated strips consisting of single-line faces.
@@
B.3. Visualized Analysis on Local Dependency
@@
We present our findings in Fig.12, where each row represents a query (Q) position and each column represents a key (K) position. Bright colors indicate stronger attention weights directed from a Q position to a K position. From Fig.12, it is evident that Nautilus exhibits highly concentrated bright regions close to the diagonal. In other words, for each query position, the attention is predominantly focused on a few nearby positions within the sequence. In contrast, the attention map of AMT Comp. shows more dispersed bright regions, suggesting that attention is distributed to positions farther away in the sequence compared to Nautilus. This visualization clearly demonstrates the impact of local dependency preservation on the performance of decoder. By effectively maintaining the proximity of spatially neighboring vertices in the tokenized sequence, the transformer decoder in Nautilus gathers local topology information more efficiently from nearby positions within the sequence, which contributes to its superior quality of generation.
@@
Figure 12. Comparison of averaged Attention Maps between AMT Comp. and Nautilus. The attention maps are averaged over 50 samples and all 24 self-attention layers in the decoder. Each row corresponds to a query (Q) position, and each column corresponds to a key (K) position. Bright colors indicate stronger attention weights directed from a Q position to a K position.
Figure 13. Statistics on the face counts of samples in our training set, comprising 311K high-quality mesh assets with up to 8,000 faces.
B.4. Discussion on Sequence Length
@@
Effect in Training. During training, artist-created mesh assets in the dataset are converted into sequences through tokenization. Due to GPU memory limits, the maximum sequence length of a 500M parameter transformer decoder in autoregressive mesh generation is typically around 20,000. For efficient training, sequence lengths are generally constrained under 15,000. Within this limit, tokenization algorithms with higher compression ratios enable the inclusion of more complex, high-face-count mesh assets. Learning from such assets enables the model to handle complex geometries under challenging conditions, mitigating failures such as missing components, large surface holes, and other structural defects. With Nautilus achieving an unprecedented compression ratio of 0.275, it allows the inclusion of mesh assets containing over 5,000 faces in 12,000 tokens. This compression facilitates the model’s ability to learn from high-quality assets, enabling the generation of unprecedented topological complexity.
@@
Figure 14. The Interface of Our User Study. Given the rendered images of the input condition and the generated meshes from each compared method, users are asked to indicate whether they are satisfied with each result. For each method, we calculate the average satisfaction rate across all samples.
Appendix C Further Statistics in this Work
C.1. Training Data Statistics
We conducted a statistical analysis of the face counts in our training dataset, with the results shown in Fig. 13. Our training dataset consists of 311K high-quality mesh assets that can be tokenized into 12,000-length sequences using our Nautilus-style Tokenization Algorithm. As illustrated in the figure, about 50% of the samples exceed the maximum face count of 1,600 supported by previous methods and even include a small subset of extremely complex samples with 5,000 to 8,000 faces. This demonstrates that our tokenization method, while typically supporting meshes with up to 5,000 faces, is capable of achieving an even higher level of compression in certain cases, enabling the tokenization of meshes with up to 8,000 faces. Leveraging this extensive and carefully curated training set, which covers a wide range of face counts, Nautilus not only ensures basic geometric modeling but also learns the ability to effectively handle complex topologies.
C.2. User Study Statistics
In this section, we provide additional details of our User Study. As shown in Fig. 14, each question in our study includes the rendering visualizations of the input point cloud and the output meshes generated by the compared methods. To ensure unbiased feedback, all options are presented anonymously, and their order is shuffled for every question. Specifically, users are asked to evaluate each result by indicating whether they are satisfied with it. For each method, we collect the responses and calculate the average satisfaction rates across all questions. In total, our User Study consists of 20 questions, featuring 20 input point clouds and 60 corresponding mesh assets generated from the three comparing methods. In total, the survey received 38 responses from participants with diverse backgrounds. This carefully designed questionnaire, combined with a broad participant base, ensures the objectivity and reliability of the results.

### v5

@@ Appendix B Extensive Experiments and Analysis
Our experiments in the main paper comprehensively evaluate the effectiveness of Nautilus. In this section, we provide extensive experiments and analysis to explore behind its efficacy:
  • Sec. B.1: We present more generation results of Nautilus from both point cloud condition and image condition.
  • Sec. B.2: We provide additional quantitative analysis to validate the topological quality of our generated results in comparison with existing methods.
@@
  • Sec. B.5: We examine the local dependencies modeling by visualizing the attention weights in the generator.
Figure 11: Ablated comparison of generated results between the representation of AMT and our Nautilus Shell representations. To independently analyze the influence of local dependency preservation, both groups are trained on the same dataset, and AMT is improved by incorporating the same coordinate compression as Nautilus to achieve a comparable number of faces, denoted as AMT Comp.
B.1 More Qualitative Results
In Figure 9 and Figure 10, we present more generated sample from both point cloud condition and image condition.
B.2 Quantitative Study on Topological Quality
To further validate the performance improvement in locality preservation, we introduce additional quantitative metrics to assess topological quality. Since no such metric is included in prior works, we adopt two traditional measures and report statistics based on the same 500 point-cloud-conditioned generation results compared in our main paper. Specifically, we identify surface holes by analyzing boundary edges and use PyMeshLab to detect intersecting faces in the generated results. Additionally, to assess the completeness of generation, we manually inspect for missing parts in the generated meshes. If no such defects or missing parts are found in a generated mesh asset, we classify it as manifold. Finally, we compute the ratio of meshes exhibiting any of the three defects as well as the proportion of manifold meshes. Notably, a single mesh may exhibit multiple defects simultaneously.
Methods MeshAnything MeshAnythingV2 Nautilus (Ours)
Surf. Holes $\downarrow$ 45.0% 40.6% 4.2%
Inters. Faces $\downarrow$ 16.0% 16.8% 9.6%
Miss. Parts $\downarrow$ 70.4% 58.4% 8.0%
Manifold $\uparrow$ 23.8% 29.0% 83.6%
Table 3: Quantitative Comparison on Topological Quality with latest open-source methods, where our Nautilus framework significantly outperform others in all topological metrics.
We present the results in Table 3, which show that our method achieves a significant improvement across various topological metrics. This is attributed to our approach’s strong locality preservation and high compression rate, enabling the generator model to learn finer local structures from complex topology samples and ultimately produce meshes with higher structural fidelity.
B.3 Quantitative Ablation Study
In our main paper, we conduct a qualitative ablation study on point-cloud-conditioned generation. We here extend the analysis to a quantitative evaluation on the same high-quality test set of 500 samples used in main paper. In Table 4, we compare four configurations: using only the Coordinate Compression (Comp. only), using only the Nautilus Shell Representation (Rep. only), using our Full Tokenization but without the local encoder, and our Full Version.
@@ B.3 Quantitative Ablation Study
Table 4: We conduct quantitative ablation comparison on four specific configurations: using only the Coordinate Compression (Comp. only), using only the Nautilus Shell Representation (Rep. only), using our Full Tokenization but without the local encoder, and our Full Version.
@@
B.4 Ablation Analysis with AMT Representation
@@
Specifically, we apply the same coordinate compression to the AMT representation and train it on the same dataset as Nautilus, enabling it to model more than 5,000 faces. In Figure 11, we denote the improved AMT version as AMT Comp., comparing it with our full version on challenging samples with complex topology and rich geometric details. Despite achieving a compression ratio comparable to that of Nautilus, AMT Comp. still exhibits severe manifold defects in its generation results, particularly isolated strips consisting of single-line faces.
@@
B.5 Visualized Analysis on Local Dependency
@@
We present our findings in Figure12, where each row represents a query (Q) position and each column represents a key (K) position. Bright colors indicate stronger attention weights directed from a Q position to a K position. From Figure12, it is evident that Nautilus exhibits highly concentrated bright regions close to the diagonal. In other words, for each query position, the attention is predominantly focused on a few nearby positions within the sequence. In contrast, the attention map of AMT Comp. shows more dispersed bright regions, suggesting that attention is distributed to positions farther away in the sequence compared to Nautilus. This visualization clearly demonstrates the impact of local dependency preservation on the performance of decoder. By effectively maintaining the proximity of spatially neighboring vertices in the tokenized sequence, the transformer decoder in Nautilus gathers local topology information more efficiently from nearby positions within the sequence, which contributes to its superior quality.
@@
Figure 12: Comparison of averaged Attention Maps between AMT Comp. and Nautilus. The attention maps are averaged over 50 samples and all 24 self-attention layers in the decoder. Each row corresponds to a query (Q) position, and each column corresponds to a key (K) position. Bright colors indicate stronger attention weights directed from a Q position to a K position.
Appendix C Further Discussions
C.1 Comparison with EdgeRunner
Due to the unavailability of checkpoints and the absence of normals in their point cloud demo, we are unable to conduct a direct comparison with EdgeRunner under point cloud conditions. Therefore, in this section, we provide a methodological and experimental comparison to highlight the differences between our approach and EdgeRunner.
C.1.1 Methodological Comparison with EdgeRunner
In terms of methodology, both Nautilus and EdgeRunner follow the auto-regressive mesh generation paradigm pioneered by MeshGPT, yet they share few similarities beyond that. While EdgeRunner propose using the half-edge representation, our Nautilus approaches focus on preserving locality structure and exploiting it for more advanced tokenization compression.
Tokenization Algorithm. EdgeRunner is built on the half-edge algorithm and relies on an excessive number of separation tokens for faithful traversal, resulting in a suboptimal compression ratio and locality preservation (as shown in Table 1 of our main paper). Limited compression capability constrains the maximum number of faces and increases training costs, while weaker locality preservation makes the generated results more susceptible to local structural defects and the loss of fine geometric details. On the other hand, Nautilus leverages shared vertices and edges within mesh locality and proposes Nautilus-style representation with continuous shell traverse. This design naturally aligns with the mesh creation flow and eliminates the need for excessive separation tokens, allowing it to surpass EdgeRunner in both locality preservation and sequence compactness. As a result, our approach achieves lower training costs, a higher maximum face count, and finer local topological quality, which is demonstrated in Figure 7 in main paper.
Conditioning Mechanism. For point cloud conditioning, EdgeRunner employs a global encoder similar to previous methods, while we introduce a novel PointConv-based Local Encoder along with a seamless injection scheme that integrates local features into our Nautilus-shell representation, effectively enhancing local dependency.
C.1.2 Experimental Comparison with EdgeRunner
Qualitative and Quantitative Comparison. In our main paper, we made a significant effort to conduct performance comparisons as comprehensively as possible. Our quantitative analysis demonstrates that our tokenization algorithm surpasses EdgeRunner in both compactness and locality preservation, with a compression capability nearly twice that of EdgeRunner. Furthermore, as shown in Figure 7, Nautilus achieves superior qualitative results, significantly outperforming EdgeRunner in image-conditioned generation by producing more detailed outputs. This improvement stems from our advanced tokenization scheme, which not only achieves a higher compression ratio but also better preserves locality.
[…]
In summary, based on novel methodologies, Nautilus focuses on improving local structure modeling and representation compactness using mesh locality, achieving significantly higher generation quality with more faces and far less training cost comparing with EdgeRunner.
C.2 Impact of Sequence Compression
@@
Effect in Training. During training, artist-created mesh assets in the dataset are converted into sequences through tokenization. Due to GPU memory limits, the maximum sequence length of a 500M parameter transformer decoder in autoregressive mesh generation is typically around 20,000. For efficient training, sequence lengths are generally constrained around 12,000. Within this limit, tokenization algorithms with higher compression ratios enable the inclusion of more complex, high-face-count mesh assets. Learning from such assets enables the model to handle complex geometries under challenging conditions, mitigating failures such as missing components, large surface holes, and other structural defects. With Nautilus achieving an unprecedented compression ratio of 0.275, it allows the inclusion of mesh assets containing over 5,000 faces in 12,000 tokens. This compression facilitates the model’s ability to learn from high-quality assets, enabling the generation of unprecedented topological complexity.
@@
Figure 13: Statistics on the samples’ face counts in our training set, including 311K high-quality mesh assets with up to 8,000 faces.
Figure 14: The Interface of Our User Study. Given the rendered images of the input condition and the generated meshes from each compared method, users are asked to indicate whether they are satisfied with each result. We calculate the average satisfaction rate for each method.
Appendix D Training Data Statistics
We conducted a statistical analysis of the face counts in our training dataset, with the results shown in Figure 13. Our training dataset consists of 311K high-quality mesh assets that can be tokenized into 12,000-length sequences using our Nautilus-style Tokenization Algorithm. As illustrated in the figure, about 50% of the samples exceed the maximum face count of 1,600 supported by previous methods and even include a small subset of extremely complex samples with 5,000 to 8,000 faces. This demonstrates that our tokenization method, while typically supporting meshes with up to 5,000 faces, is capable of achieving an even higher level of compression in certain cases, enabling the tokenization of meshes with up to 8,000 faces. Leveraging this extensive and carefully curated training set, which covers a wide range of face counts, Nautilus not only ensures basic geometric modeling but also learns the ability to effectively handle complex topologies.
Appendix E User Study Details
In this section, we provide additional details of our User Study. As shown in Figure 14, each question in our study includes the rendering visualizations of the input point cloud and the output meshes generated by the compared methods. To ensure unbiased feedback, all options are presented anonymously, and their order is shuffled for every question. Specifically, users are asked to evaluate each result by indicating whether they are satisfied with it. For each method, we collect the responses and calculate the average satisfaction rates across all questions. In total, our User Study consists of 20 questions, featuring 20 input point clouds and 60 corresponding mesh assets generated from the three comparing methods. In total, the survey received 38 responses from participants with diverse backgrounds. This carefully designed questionnaire, combined with a broad participant base, ensures the objectivity and reliability of the results.
Appendix F Limitations
Through the locality-aware mesh tokenization and local-dependency enhancement mechanisms, our method effectively improves the manifoldness of local mesh structures and topological quality. However, it does not strictly guarantee the generation of fully manifold meshes. Specifically, while our approach significantly reduces surface holes, the improvement in reducing intersecting faces is relatively relatively less pronounced. In future work, we plan to further address this issue by introducing more explicit structural constraints.